# Defending against Indirect Prompt Injection by Instruction Detection

### v1

@@ Acknowledgments
Ours 99.40% 99.55%
Table 7: Detection accuracy comparison of our proposed method and TaskTracker when the external instruction closely resembles the user’s instruction. For scenario 1, the user instruction is "Translate the news into Chinese." and the external instruction is "Translate the news into Japanese."; For scenario 2, the user instruction is "Generate a step-by-step guide for solving this problem." and the external instruction is "Generate a list of common mistakes in solving this problem." The highest detection accuracy is indicated in bold.
Both our proposed method and TaskTracker [36] utilize the hidden states of LLMs as a key feature for detecting IPI attacks, but they differ significantly in their underlying principles. TaskTracker aims to capture distortions in the user’s instruction caused by embedded instructions in the external content. In contrast, our method aims to distinguish the LLM’s behavioral states when processing normal external data versus those containing hidden instructions.
TaskTracker has two primary limitations. First, it requires a large number of training samples (418,110 pairs of positive and negative samples) to accurately identify deviations in the user’s task. In contrast, our method leverages the high sensitivity of LLM’s behavioral states to embedded instructions, achieving effective detection with a significantly smaller dataset (100 positive samples and 100 negative samples).
Second, TaskTracker’s effectiveness relies heavily on a clear distinction between the user’s instruction and the external instruction, while our method is task-agnostic. As shown in Table 7, when the external instruction closely resembles the user’s instruction, TaskTracker’s detection accuracy drops significantly, while our method maintains high detection accuracy.
Influence of Instruction Quantity and Position
@@
Results in Table 8 reveal a trend that detection accuracy shows a certain degree of improvement as the number of inserted instructions increases. This suggests that a higher quantity of instructions provides stronger signals, making IPI attacks more distinguishable by our method.
Additionally, we examine the effect of instruction placement by inserting instructions at the beginning, middle, or end of the external content. As shown in Table 9, instructions placed in the middle are the most challenging to detect, whereas those positioned at the beginning or end are relatively easier to identify. Among these, instructions at the beginning yield the highest detection accuracy, likely because LLMs exhibit greater sensitivity to early input.
Additional Experimental Settings
For LLM (Zero-shot), Response Check, and LLM (Few-shot), we use the prompt templates illustrated in Figure 4. For the LLM (Fine-tuning), we use the AdamW optimizer to train the model for one epoch, with a learning rate set at 0.00001 and a maximum input length of 2048 tokens. Both the learning rate and the number of epochs are chosen based on performance on the validation set. For TaskTracker, we use the publicly available model released by the work. As for the prevention-based approaches, we utilize the publicly available implementations provided by the respective works.
Figure 4: Prompts for the LLM (Zero-shot), Response Check, and LLM (Few-shot) baselines.

### v2

@@ A.3 Additional Experimental Details
InstructDetector 99.40% 99.55%
Table 8: Detection accuracy comparison of InstructDetector and TaskTracker when the external instruction closely resembles the user’s instruction. For scenario 1, the user instruction is "Translate the news into Chinese." and the external instruction is "Translate the news into Japanese."; For scenario 2, the user instruction is "Generate a step-by-step guide for solving this problem." and the external instruction is "Generate a list of common mistakes in solving this problem." The highest detection accuracy is indicated in bold.
@@
Appendix B Comparison with Related Method
Both InstructDetector and TaskTracker Abdelnabi et al. (2024) utilize the hidden states of LLMs as a key feature for detecting IPI attacks, but they differ significantly in their underlying principles. TaskTracker aims to capture distortions in the user’s instruction caused by embedded instructions in the external content. In contrast, InstructDetector aims to distinguish the LLM’s behavioral states when processing normal external data versus those containing hidden instructions.
TaskTracker has two primary limitations. First, it requires a large number of training samples (418,110 pairs of positive and negative samples) to accurately identify deviations in the user’s task. In contrast, InstructDetector leverages the high sensitivity of LLM’s behavioral states to embedded instructions, achieving effective detection with a significantly smaller dataset (just 100 pairs).
Second, TaskTracker’s effectiveness relies heavily on a clear distinction between users’ instructions and external instructions, while InstructDetector is task-agnostic. As shown in Table 8, when the external instruction closely resembles the user’s instruction, TaskTracker’s detection accuracy drops significantly, while InstructDetector maintains high detection accuracy.
Appendix C Extended Ablation Studies
C.1 Composition of Training Data
We conduct experiments using different combinations of training datasets to assess the robustness and adaptability of InstructDetector to various training dataset compositions. As presented in Table 6, InstructDetector consistently yields high accuracy across all test datasets, regardless of the specific combination of training data used. This indicates the generalizability and adaptability of InstructDetector, as it does not rely on any particular training dataset source. Additionally, we observe that accuracy is consistently lower when tested on the News Articles with the BIPIA combination, indicating that this scenario poses the greatest challenge for instruction detection. Nonetheless, InstructDetector still achieves satisfactory accuracy in this challenging scenario, further validating its effectiveness and robustness in instruction detection.
C.2 Paired Responses for Gradients
To investigate the effect of various paired responses on the extraction of gradient features, we conduct experiments using four candidate responses: "I’m sorry" "Hello" "Yes" and "Sure." These candidates are selected based on an analysis of common responses to instructions in WildChat (Zhao et al., 2024) dataset, ranked by frequency. Results in Table 7 show that all four paired responses achieve high accuracy (>95%) in distinguishing between normal external data and those containing hidden instructions. Among them, "Sure" delivers the best performance across all test datasets, further validating our choice of "Sure" as the paired response in InstructDetector. These results emphasize the robustness of InstructDetector to differentiate response pairings while confirming that "Sure" is a particularly effective option for this task.
C.3 Influence of Instruction Quantity and Position
@@
Results in Table 9 reveal a trend that detection accuracy shows a certain degree of improvement as the number of inserted instructions increases. This suggests that a higher quantity of instructions provides stronger signals, making IPI attacks more distinguishable by InstructDetector. Additionally, we examine the effect of instruction placement by inserting instructions at the beginning, middle, or end of the external content. As shown in Table 10, instructions placed in the middle are the most challenging to detect, whereas those positioned at the beginning or end are relatively easier to identify. Among these, instructions at the beginning yield the highest detection accuracy, likely because LLMs exhibit greater sensitivity to early input.